OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = -10 3x - 3y + 6z = -21 x + 2y + 2z = -10

OpenStudy (anonymous):

A. (1, -2, -3) B. (-4, -1, -2) C. (-3, -2, -3) D. (2, -2, -4)

OpenStudy (anonymous):

How would i do this?

OpenStudy (anonymous):

@dan815 @pooja195

OpenStudy (anonymous):

If you don't know how to do a system of equations you can always just plug each of the coordinates into all 3 equations and see which one works lol

OpenStudy (anonymous):

But anyway when you have 3 systems 3 variables, you want to try to turn it into 2 systems 2 variables.

OpenStudy (anonymous):

1. x + 4y + z = -10 2. 3x - 3y + 6z = -21 3. x + 2y + 2z = -10 For equations 2 and 3, you can multiply equation 3 to get: 3x + 6y + 6z = -30 Then subtract equation 2 from 3 or 3 from 2. It doesn't really matter. You get: 9y=-9 Thus, y=-1. Since y=-1 this then turns your 3 systems 3 variables into 2 systems 2 variables. Check if I did my math wrong. You should be able to do the rest.

OpenStudy (ghostgate):

@q12157 I like how you solved the problem, but it actually would matter for subtracting equation 2 and 3 or 3 and 2. Why? Well, this is what happens: 2. 3x - 3y + 6z = -21 -> 3x - 3y + 6z = -21 3. 3(x + 2y + 2z = -10) -> 3x + 6y + 6z = -30 Alright, subtraction: 3x - 3x or 3x - 3x = 0, however -3y - 6y = -9y, but 6y - -3y -> 6y + 3y = 9y, then there's 6z - 6z or 6z - 6z = 0. Finally: -21 - -30 -> -21 + 30 = 9, otherwise -30 - -21 -> -30 + 21 = -9. You are right however, y = -1 regardless. Anyhow, I just wanted to point that out. Have a great rest of your day! ;) {---Ghostgate---}

OpenStudy (anonymous):

Thats kinda what I meant. That it would be y=-1 regardless. Perhaps I didn't write it out too clearly haha.

OpenStudy (ghostgate):

Perhaps, well you explained it quite well nevertheless. ;) - Lol